# Mind the Gap: The Divergence Between Human and LLM-Generated Tasks

### v2

@@ Discussion
This study examined the cognitive basis of autonomous goal generation by comparing human behavior with large language models (LLMs). Our findings show that human-generated goals are systematically shaped by personal values and are adapted to environmental complexity. In contrast, LLMs, even when prompted with personal profiles, exhibit fundamental and predictable biases, revealing a significant gap between their text-based simulation and the grounded nature of human cognition.
@@
Individual differences shaped goal content. People high in openness to change generated more novel and fun tasks, consistent with prior work linking openness to exploratory behavior and intrinsic motivation (Deci and Ryan 2013; McCrae and Costa Jr 1997; Sagiv and Roccas 2021). Cognitive style also mattered: systematic thinkers generated less diverse tasks in complex environments, possibly due to higher cognitive load. These results align with previous creativity and decision-making theories, which suggest that systematic thinkers prefer structured problem-solving, while intuitive thinkers are more adaptable in diverse settings (Evans and Stanovich 2013; Sagiv et al. 2014).
@@
First is the embodiment gap. Humans generated tasks involving physical interaction with common objects (e.g., hangers, pillows), reflecting sensorimotor knowledge of object use. LLMs rarely produced such tasks. Their outputs were rated as less physically demanding and less involving of limbs. This supports the view that LLMs rely on text-based associations and lack access to intuitive physics or affordance perception (Lake et al. 2017).
Second is the value and motivation gap. Humans often proposed social and prosocial tasks, consistent with basic motivational drives such as affiliation and well-being. In contrast, LLMs showed a lower propensity for social engagement. Their strong bias toward abstract, mental tasks (e.g., music, writing) and away from relaxation or household chores points to a lack of an internal, value-driven system. Crucially, providing the model with user’s psychological profile did not fix this. This finding suggests that simply describing values in a prompt is not equivalent to having an internal value systems that regulates task selection.
@@
LLM-generated tasks were rated as more novel and fun. This highlights the unique nature of LLM creativity. Unconstrained by physical feasibility, LLMs excel at combinatorial creativity, drawing on their vast training data to produce imaginative textual descriptions.
This natural tendency is likely amplified by current alignment methods like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022). Such methods optimize for outputs that human raters find immediately helpful, engaging, or interesting. This process inherently discourages the generation of mundane, routine, or repetitive actions like resting or cleaning. Although these activities are central to daily life, they are less likely to receive high ratings and are therefore filtered out. The result is a model fine-tuned to produce “ungrounded” but highly-rated novelties, which explains both its high creativity scores and its thematic detachment from the reality of human experience.
@@
Our findings have two main implications. For AI research, they suggest that creating truly human-like autonomous agents requires more than scaling up language models. Addressing the embodiment and motivation gaps may require integrating world models, intrinsic reward systems, or sensorimotor learning (Matsuo et al. 2022). For cognitive science, our task-based paradigm offers a scalable, flexible tool to study goal-setting in naturalistic contexts, beyond traditional lab constraints. It enables new comparisons between human cognition and machine outputs.
Several limitations of our study point toward promising future directions. First, the text-based task generation paradigm may not fully capture the richness of human goal-setting, which often involves multimodal, interactive, and context-sensitive elements. Future work could incorporate interactive 3D environments to enhance ecological validity. Second, due to the high cost of human evaluation, we restricted our comparisons to a single LLM architecture (GPT-4o). Expanding the scope to include a broader range of models with different sizes and architectures could offer a more comprehensive understanding of how different model properties influence goal generation. Finally, the diversity and complexity of human-generated goals highlight the limitations of current computational frameworks, and call for new approaches capable of modeling open-ended, value-guided behavior.
Acknowledgments

### v3

@@ Discussion
This study examined the cognitive basis of autonomous goal generation by comparing human behavior with LLM. Our findings show that human-generated goals are systematically shaped by personal values and adapted to environmental complexity. In contrast, LLMs, even when prompted with personal profiles, exhibit a fundamental and predictable gap between their text-based simulation and the grounded nature of human cognition.
@@
Individual differences shaped goal content. People high in openness to change generated more novel and fun tasks, consistent with prior work linking openness to exploratory behavior and intrinsic motivation (Deci and Ryan 2013; McCrae and Costa Jr 1997; Sagiv and Roccas 2021). Cognitive style also mattered: systematic thinkers generated less diverse tasks in complex environments, possibly due to higher cognitive load. These results align with theories of creativity and decision-making, which suggest that systematic thinkers prefer structured problem-solving, while intuitive thinkers are more adaptable in diverse settings (Evans and Stanovich 2013; Sagiv et al. 2014).
@@
First is the embodiment gap. Humans generated tasks involving physical interaction with common objects (e.g., hangers, pillows), reflecting sensorimotor knowledge of object use. LLMs rarely produced such tasks, and their outputs were rated as less physically demanding and less involving of limbs. This supports the view that LLMs rely on text-based associations and lack access to intuitive physics or affordance perception (Lake et al. 2017).
Second is the value and motivation gap. Humans often proposed social and prosocial tasks, consistent with basic motivational drives like affiliation. In contrast, LLMs showed a lower propensity for social engagement and a strong bias toward abstract, mental tasks (e.g., music, writing). Crucially, providing the model with a user’s psychological profile did not fix this. This finding suggests that describing values in a prompt is no substitute for an integrated value system that actively regulates goal selection.
@@
LLM-generated tasks were rated as more novel and fun. This highlights the unique nature of LLM creativity. Unconstrained by physical feasibility, LLM excel at combinatorial creativity, drawing on their vast training data to produce imaginative textual descriptions.
This natural tendency is likely amplified by current alignment methods like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022). Such methods optimize for outputs that human raters perceive as immediately helpful or engaging. This reveals a fundamental tension between two objectives: developing an expert LLM skilled in creative or intellectual domains versus constructing one that mirrors the often mundane patterns of real human behavior. This optimization paradigm inherently discourages the generation of routine actions like resting or cleaning. The result is a model fine-tuned to produce “ungrounded” but highly-rated novelties, explaining both its high creativity scores and its thematic detachment from the reality of human experience.
@@
Our findings have two main implications. For AI research, they suggest that creating truly human-like autonomous agents requires more than scaling up LLM. Addressing the embodiment and motivation gaps may require integrating world models, intrinsic reward systems, or sensorimotor learning (Matsuo et al. 2022). For cognitive science, our task-based paradigm offers a scalable, flexible tool to study goal-setting in naturalistic contexts, beyond traditional lab constraints. It enables new comparisons between human cognition and machine outputs.
Several limitations of our study point toward promising future directions. First, the text-based task generation paradigm may not fully capture the richness of human goal-setting, which often involves multimodal, interactive, and context-sensitive elements. Future work could incorporate interactive 3D environments with more scenarios to enhance ecological validity. Second, due to the high cost of human evaluation, our study focused on a deep, theory-driven analysis of a single architecture (GPT-4o), which represented the state-of-the-art at the time of our pre-registration. This approach prioritizes depth over breadth, establishing a detailed cognitive baseline. Future work should expand these comparisons across a broader range of models to test the generality of our findings. Finally, the complexity of human-generated goals calls for new computational frameworks capable of modeling open-ended, value-guided behavior.
Acknowledgements